OpenStudy (anonymous):

HELP: how do you find the smallest number divisible by both 715 and 770?

OpenStudy (rmrjr22):

Divide by 5?

OpenStudy (dumbcow):

yeah its 5

OpenStudy (anonymous):

the answer is 10010

OpenStudy (anonymous):

its the LCM you're looking for 770=2x5x7x11 715=5x11x13 so LCM(770;715)=2x5x7x11x13=10010

OpenStudy (anonymous):

how do you go about finding the LCM?

OpenStudy (anonymous):

factor each number 715 = 5 * 11 * 13 as ghass says

OpenStudy (anonymous):

first do the prime factorization of each number the LCM is the product of all factors raised to the highest power

OpenStudy (anonymous):

how do you know which are raised to the highest power ?

OpenStudy (anonymous):

well in our case all the powers are 1 but for example if a=2^2x3^4 and b=2x3^2x7 then LCM(a,b)=2^2x3^4x7=...